Anna has just received a gift of $500 for her graduation, which increased her net worth by $500. If she uses the money to purchase a video player how will her net worth be affected? If she invests the $500 at 10% interest per year, what will it be worth in one year?
Answer by ikleyn(52781)

.

I will answer the second question, ONLY.

    It will be worth  500 + 0.10*500 = 500 + 50 = 550  dollars.    ANSWER

Solved.